OpenStudy (anonymous):

solve for y in the following system of equations x-y=-1 3x+5y=21

OpenStudy (anonymous):

x = y-1 3y -3 +5y =21 8y=24 y=3 x = 2

OpenStudy (anonymous):

could you please explain the first two steps briefly?

OpenStudy (anonymous):

x - y = -1 x = y -1 add y both sides then 3(x) +5y = 21 3(y-1)+5y =21